# Learning to Translate Noise for Robust Image Denoising

### v1

@@ B.4 Expanded Table: Gaussian Injection Block and Explicit Noise Translation Loss
Table S.5: Quantitative comparisons of real-world denoising performance across various training set configurations. Methods marked with an asterisk (*) are evaluated using official out-of-the-box models.
@@
Denoising network Trained noise Metric SIDD   Poly    CC HighISO iPhone Huawei OPPO Sony Xiaomi Total Avg.
   Restormer [36] SIDD* PSNR 39.93 37.63 36.31 38.24 40.05 38.36 39.49 44.02 35.62 38.85
SSIM 0.9598 0.9790 0.9805 0.9753 0.9727 0.9671 0.9768 0.9889 0.9706 0.9745
@@
Ours PSNR 39.08 38.74 37.60 40.06 41.62 39.68 40.55 44.12 36.14 39.73
SSIM 0.9558 0.9846 0.9861 0.9851 0.9751 0.9761 0.9794 0.9849 0.9747 0.9779
   NAFNet [6] SIDD* PSNR 39.88 37.13 35.81 38.27 40.16 37.68 39.57 43.48 34.97 38.55
SSIM 0.9594 0.9714 0.9809 0.9784 0.9700 0.9676 0.9781 0.9824 0.9681 0.9729
@@
Ours PSNR 39.17 38.67 37.82 39.94 41.94 39.74 40.45 44.17 36.14 39.78
SSIM 0.9566 0.9851 0.9876 0.9853 0.9805 0.9778 0.9796 0.9869 0.9745 0.9793
Table S.6: Quantitative results of applying noise translation framework to denosing model pre-trained on pure Gaussian.
Denoising methods Metric SIDD   Poly    CC HighISO iPhone Huawei OPPO Sony Xiaomi Total Avg.
@@
B.5 Comparison with naïve combination of real and synthetic dataset for training
We compare our method with the naive approach of combining Gaussian noise dataset with the SIDD dataset for training. The results are presented in Table S.5. The Gaussian noise dataset used in this comparison is identical to the one which is used for our method. The results demonstrate that naively combining real-noise and Gaussian noise datasets during training does not yield significant improvements in generalization performance. This highlights the superiority and efficacy of our method in enhancing denoising performance across diverse noise distributions.
B.6 Incorporating a pure Gaussian pretrained denoising network
We conduct an experiment on the officially published denoising model, Restormer [36], which was trained for blind gaussian noise removal. By leveraging it in our framework, we trained the corresponding noise translation network with the proposed methodology to evaluate the overall performance. The results are provided in Table S.6, which demonstrate that even when using a pure Gaussian pretrained out-of-the-box model within our framework, the real-world denoising performance undergoes significant improvement. This enhancement can be attributed to our noise translation network, which effectively adapts the real-world noise into a form closer to the Gaussian noise that the pretrained model is adept at handling. Note that, in our pretraining approach, the denoising network is trained on Gaussian noise augmented with the same real-world noise used to train the noise translation network. This distinction in pretraining strategies explains the observed performance gap between the two results.
B.7 Effects of hyperparameters
Table S.7: Sensitivity results on various hyperparameters in our framework.
@@ B.7 Effects of hyperparameters
5 39.37 38.96 0 39.37 38.66 0 39.05 39.61 0 39.10 39.78
10 39.29 39.64 1 39.33 39.24 0.001 39.03 39.65 0.001 39.15 39.82
15 39.17 39.86 5 39.17 39.66 0.005 39.07 39.72 0.002 39.17 39.86
20 38.88 39.53 15 39.08 39.77 0.01 39.09 39.76 0.005 39.07 39.82
25 38.73 39.21 50 39.11 39.80 0.05 39.17 39.86 0.01 39.07 39.56
100 39.17 39.86 0.1 39.05 39.67 0.02 39.07 38.97
200 39.05 39.70 0.5 38.90 37.80 0.05 38.57 37.41
Table S.7 presents the ablative results of our hyperparameters, including pretraining noise level ($\sigma$), noise injection level ($\tilde{\sigma}$), explicit loss weight ($\alpha$), and spatial frequency ratio ($\beta$). The pretraining level part of Table S.7 shows the ablation results for the Gaussian noise levels added to create noisy input during denoising network pretraining. As the noise level increased, the performance on in-distribution (ID) consistently decreased. For out-of-distribution (OOD), the performance improved until noise level of 15, beyond which it began to degrade, due to oversmoothing effects caused by learning to handle strong noise. The noise injection level part presents the ablation results for Gaussian noise injection levels. Increasing the noise level led to a decline in ID performance, while OOD performance improved up to a certain point. The explicit loss weight and spatial-frequency ratio parts show the ablation results for the $\alpha$ and $\beta$ values in Eq. (12) and (10), respectively. Overall, the ablation experiments determined the optimal hyperparameters as follows: a pretraining noise level $\sigma=15$, a Gaussian noise injection level $\tilde{\sigma}=100$, $\alpha=5\times 10^{-2}$, and $\beta=2\times 10^{-3}$.
B.8 Training Stability
@@
Refer to caption iPhone Noisy Refer to caption AP-BSN 42.51 dB Refer to caption SSID 43.80 dB Refer to caption Restormer 41.73 dB Refer to caption Ground Truth
Refer to caption NAFNet 42.52 dB Refer to caption KBNet 42.25 dB Refer to caption Xformer 41.71 dB Refer to caption Ours 46.92 dB
Refer to caption Huawei Noisy Refer to caption AP-BSN 27.67 dB Refer to caption SSID 29.47 dB Refer to caption Restormer 33.75 dB Refer to caption Ground Truth
Refer to caption NAFNet 31.13 dB Refer to caption KBNet 27.98 dB Refer to caption Xformer 33.58 dB Refer to caption Ours 34.05 dB
Refer to caption OPPO Noisy Refer to caption AP-BSN 40.30 dB Refer to caption SSID 40.02 dB Refer to caption Restormer 39.37 dB Refer to caption Ground Truth
Refer to caption NAFNet 38.90 dB Refer to caption KBNet 31.03 dB Refer to caption Xformer 39.09 dB Refer to caption Ours 42.08 dB
Figure S.1: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet) on iPhone, Huawei, and OPPO dataset. Zoom in for a closer inspection of the visual quality.
Refer to caption Sony Noisy Refer to caption AP-BSN 39.43 dB Refer to caption SSID 40.13 dB Refer to caption Restormer 41.53 dB Refer to caption Ground Truth
Refer to caption NAFNet 40.71 dB Refer to caption KBNet 40.94 dB Refer to caption Xformer 41.40 dB Refer to caption Ours 41.68 dB
Refer to caption Xiaomi Noisy Refer to caption AP-BSN 34.68 dB Refer to caption SSID 36.91 dB Refer to caption Restormer 38.39 dB Refer to caption Ground Truth
Refer to caption NAFNet 36.09 dB Refer to caption KBNet 36.85 dB Refer to caption Xformer 38.29 dB Refer to caption Ours 39.45 dB
Refer to caption SIDD Noisy Refer to caption AP-BSN 30.63 dB Refer to caption SSID 31.83 dB Refer to caption Restormer 34.81 dB Refer to caption Ground Truth
Refer to caption NAFNet 34.33 dB Refer to caption KBNet 34.31 dB Refer to caption Xformer 34.88 dB Refer to caption Ours 34.11 dB
Figure S.2: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet) on Sony, Xiaomi, and SIDD datasets. Zoom in for a closer inspection of the visual quality.
Refer to caption High ISO Refer to caption Desnoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Desnoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Desnoised High ISO (Ours) Refer to caption ISO 200
Figure S.3: Qualitative results of our noise translation network with pretrained NAFNet applied to high ISO images from the NIND dataset. Our method effectively removes noise while preserving fine details, producing outputs comparable to low ISO (ISO 200) images. Zoom in for a closer inspection of the visual quality.
Refer to caption High ISO Refer to caption Desnoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Desnoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Desnoised High ISO (Ours) Refer to caption ISO 200
Figure S.4: Qualitative results of our noise translation network with pretrained NAFNet applied to high ISO images from the NIND dataset. Our method effectively removes noise while preserving fine details, producing outputs comparable to low ISO (ISO 200) images. Zoom in for a closer inspection of the visual quality.
C Qualitative Analysis
C.1 Additional OOD datasets and SIDD
@@
Qualitative results on all datasets in the main paper are shown in Figures S.1 and S.2. Our method significantly surpasses the PSNR scores of other denoising models on all OOD datasets (iPhone, Huawei, OPPO, Sony, Xiaomi). Notably, the output images of KBNet [42], exhibit visible breakdowns due to severe overfitting to the training set. For the in-distribution SIDD, our method produces clean results without generating unnecessary zipper artifacts.
C.2 NIND (Natural Image Noise Dataset)
@@
The NIND [3] dataset comprises high-resolution images captured at various ISO levels. High ISO images are noisier due to increased sensor sensitivity, while low ISO images, though cleaner, may require more optimal settings or equipment to capture effectively. Figure S.3 and Figure S.4 illustrate the qualitative results of our method applied to high ISO images, demonstrating its ability to effectively remove noise while preserving fine image details. The denoised outputs are visually clean and free from artifacts, achieving a quality comparable to images captured at low ISO levels.
C.3 Real-world smartphone image noise and denoising results
@@
Finally, we present the denoised results of images captured using our Galaxy S22+ smartphone, as shown in Figure S.5.
@@
Figure S.5: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet) on images captured by Galaxy S22+ smartphone. Zoom in for a closer inspection of the visual quality.
D Reproducibility Statement
To ensure reproducibility of our work, we have made efforts to include key details in the main text and appendix. Furthermore, we provide the complete source code and model weights in the supplementary material to facilitate easy reproduction of the experiments and results.

### v2

@@ 8.5 Expanded Table: Adaptability of the Noise Translation Network
Table 12: Quantitative comparisons of real-world denoising performance across various training set configurations. Methods marked with an asterisk (*) are evaluated using official out-of-the-box models.
@@
Denoising network Trained noise Metric SIDD   Poly    $C$ HighISO iPhone Huawei OPPO Sony Xiaomi Total Avg.
Restormer [43] SIDD* PSNR 40.02 37.66 36.33 38.29 40.13 38.42 39.56 44.19 35.65 38.92
SSIM 0.9603 0.9793 0.9807 0.9756 0.9734 0.9675 0.9773 0.9894 0.9710 0.9749
@@
Ours PSNR 39.22 38.76 37.68 40.14 41.85 39.72 40.64 44.29 36.19 39.83
SSIM 0.9569 0.9854 0.9866 0.9857 0.9786 0.9766 0.9800 0.9871 0.9750 0.9791
NAFNet [6] SIDD* PSNR 39.97 37.17 35.69 38.32 40.25 37.73 39.64 43.65 34.99 38.60
SSIM 0.9600 0.9717 0.9811 0.9788 0.9707 0.9680 0.9786 0.9829 0.9685 0.9734
@@
Ours PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.86
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9797
Table 13: Quantitative results of applying noise translation framework to denoising model pre-trained on pure synthetic Gaussian noise.
Denoising methods Metric SIDD   Poly    $C$ HighISO iPhone Huawei OPPO Sony Xiaomi Total Avg.
@@
8.6 Comparison with naïve combination of real and synthetic dataset for training
We compare our method with the naive approach of combining Gaussian noise dataset with the SIDD dataset for training. The results are presented in Table 12. The Gaussian noise dataset used in this comparison is identical to the one which is used for our method. The results demonstrate that naively combining real-noise and Gaussian noise datasets during training does not yield significant improvements in generalization performance. This highlights the superiority and efficacy of our method in enhancing denoising performance across diverse noise distributions.
8.7 Incorporating a pure Gaussian pretrained denoising network
We conduct an experiment on the officially published denoising model, Restormer [43], which was trained for blind gaussian noise removal. By leveraging it in our framework, we trained the corresponding noise translation network with the proposed methodology to evaluate the overall performance. The results are provided in Table 13, which demonstrate that even when using a pure Gaussian pretrained out-of-the-box model within our framework, the real-world denoising performance undergoes significant improvement. This enhancement can be attributed to our noise translation network, which effectively adapts the real-world noise into a form closer to the Gaussian noise that the pretrained model is adept at handling. Note that, in our pretraining approach, the denoising network is trained on Gaussian noise augmented with the same real-world noise used to train the noise translation network. This distinction in pretraining strategies explains the observed performance gap between the two results.
8.8 Effects of hyperparameters
Table 14: Sensitivity results on various hyperparameters in our framework.
@@ 8.8 Effects of hyperparameters
5 39.37 38.96 0 39.43 39.37 0 39.07 39.74 0 39.13 39.84
10 39.29 39.64 1 39.38 39.54 0.001 39.08 39.76 0.0005 39.16 39.88
15 39.24 39.94 5 39.09 39.78 0.005 39.11 39.80 0.001 39.18 39.91
20 38.88 39.53 15 39.15 39.86 0.01 39.14 39.85 0.002 39.24 39.94
25 38.73 39.21 50 39.15 39.90 0.05 39.24 39.94 0.005 39.17 39.94
100 39.24 39.94 0.1 39.19 39.92 0.01 39.09 39.77
200 39.19 39.93 0.5 39.13 39.07 0.02 39.00 38.97
Table 14 presents the ablative results of our hyperparameters, including pretraining noise level ($\sigma$), noise injection level ($\tilde{\sigma}$), explicit loss weight ($\alpha$), and spatial frequency ratio ($\beta$). The pretraining level part of Table 14 shows the ablation results for the Gaussian noise levels added to create noisy input during denoising network pretraining. As the noise level increased, the performance on in-distribution (ID) consistently decreased. For out-of-distribution (OOD), the performance improved until noise level of 15, beyond which it began to degrade, due to oversmoothing effects caused by learning to handle strong noise. The noise injection level part presents the ablation results for Gaussian noise injection levels. Increasing the noise level led to a decline in ID performance, while OOD performance improved up to a certain point. The explicit loss weight and spatial-frequency ratio parts show the ablation results for the $\alpha$ and $\beta$ values in Eq. (11) and (10), respectively. Overall, the ablation experiments determined the optimal hyperparameters as follows: a pretraining noise level $\sigma=15$, a Gaussian noise injection level $\tilde{\sigma}=100$, $\alpha=5\times 10^{-2}$, and $\beta=2\times 10^{-3}$.
8.9 Training Stability
@@
9 Qualitative Analysis
9.1 Additional OOD datasets and SIDD dataset
@@
Qualitative results on all datasets in the main paper are shown in Figures 9,  10,  11 and  12. Our method significantly surpasses the PSNR scores of other denoising models on all OOD datasets (iPhone, Huawei, OPPO, Sony, Xiaomi). For the in-distribution SIDD, our method produces clean results without generating unnecessary zipper artifacts.
9.2 NIND (Natural Image Noise Dataset)
@@
The NIND [3] dataset comprises high-resolution images captured at various ISO levels. High ISO images are noisier due to increased sensor sensitivity, while low ISO images, though cleaner, may require more optimal settings or equipment to capture effectively. Figure 13 and Figure 14 illustrate the qualitative results of our method applied to high ISO images, demonstrating its ability to effectively remove noise while preserving fine image details. The denoised outputs are visually clean and free from artifacts, achieving a quality comparable to images captured at low ISO levels.
9.3 Real-world smartphone image noise and denoising results
@@
Finally, we present the denoised results of images captured using our Galaxy S22+ smartphone, as shown in Figure 15.
10 Limitations
Although our approach provides strong robustness to diverse real-world noise, it may appear to compromise performance on in-distribution (ID) datasets. As discussed in Figure 8, this effect is largely due to the pronounced overfitting of conventional denoising models, which tend to reconstruct dataset-specific artifacts rather than genuine image content.
Moreover, our evaluation primarily relies on PSNR and SSIM computed against the provided ground-truth images. While these metrics are standard practice, they inherently struggle to capture perceptual realism and may not align well with human judgments. Although our method consistently improves these metrics, such quantitative measures do not fully represent perceptual quality, which remains subjective and can vary across viewers. Future work includes exploring alternative evaluation protocols or perceptually aligned metrics that better account for realism, semantic consistency, and human preference.
Poly Noisy
R2R 39.84dB
AP-BSN 42.13dB
SSID 42.14dB
APR-RD 42.16dB
Mask-DN 41.03dB
CLIP-DN 41.06dB
AFM 41.51dB
IDF 40.72dB
NAFNet 40.71dB
XFormer 40.35dB
Ours 43.81dB
Ground Truth
iPhone Noisy
R2R 40.94dB
AP-BSN 42.61dB
SSID 43.93dB
APR-RD 41.70dB
Mask-DN 42.60dB
CLIP-DN 42.13dB
AFM 41.85dB
IDF 41.82dB
NAFNet 42.45dB
XFormer 41.80dB
Ours 47.07dB
Ground Truth
Figure 9: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet), on the out-of-distribution (OOD) datasets (Poly and iPhone). Our result displays cleaner outputs compared to other state-of-the-art networks that are directly trained from a real-noise dataset. Zoom in for better comparison.
Huawei Noisy
R2R 33.51dB
AP-BSN 27.67dB
SSID 29.47dB
APR-RD 29.21dB
Mask-DN 31.93dB
CLIP-DN 31.45dB
AFM 32.33dB
IDF 28.33dB
NAFNet 31.92dB
XFormer 33.59dB
Ours 33.95dB
Ground Truth
OPPO Noisy
R2R 39.19dB
AP-BSN 40.36dB
SSID 40.07dB
APR-RD 40.34dB
Mask-DN 39.16dB
CLIP-DN 39.58dB
AFM 39.69dB
IDF 38.70dB
NAFNet 39.02dB
XFormer 39.14dB
Ours 42.19dB
Ground Truth
Figure 10: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet), on the out-of-distribution (OOD) datasets (Huawei and OPPO). Our result displays cleaner outputs compared to other state-of-the-art networks that are directly trained from a real-noise dataset. Zoom in for better comparison.
Sony Noisy
R2R 40.01dB
AP-BSN 39.48dB
SSID 40.19dB
APR-RD 39.58dB
Mask-DN 41.78dB
CLIP-DN 41.52dB
AFM 41.20dB
IDF 40.11dB
NAFNet 41.62dB
XFormer 41.48dB
Ours 41.72dB
Ground Truth
Xiaomi Noisy
R2R 37.99dB
AP-BSN 34.70dB
SSID 36.94dB
APR-RD 37.97dB
Mask-DN 37.91dB
CLIP-DN 38.07dB
AFM 38.76dB
IDF 34.96dB
NAFNet 37.96dB
XFormer 38.32dB
Ours 39.65dB
Ground Truth
Figure 11: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet), on the out-of-distribution (OOD) datasets (Sony and Xiaomi). Our result displays cleaner outputs compared to other state-of-the-art networks that are directly trained from a real-noise dataset. Zoom in for better comparison.
SIDD Noisy
R2R 28.98dB
AP-BSN 30.65dB
SSID 31.84dB
APR-RD 32.29dB
Mask-DN 32.92dB
CLIP-DN 30.41dB
AFM 32.19dB
IDF 31.18dB
NAFNet 34.54dB
XFormer 34.90dB
Ours 34.12dB
Ground Truth
Figure 12: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet), on the SIDD dataset. Our result displays competitive visual quality with real noise.
Refer to caption High ISO Refer to caption Denoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Denoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Denoised High ISO (Ours) Refer to caption ISO 200
Figure 13: Qualitative results of our noise translation network with pretrained NAFNet applied to high ISO images from the NIND dataset. Our method effectively removes noise while preserving fine details, producing outputs comparable to low ISO (ISO 200) images. Zoom in for a closer inspection of the visual quality.
Refer to caption High ISO Refer to caption Denoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Denoised High ISO (Ours) Refer to caption ISO 200
Refer to caption High ISO Refer to caption Denoised High ISO (Ours) Refer to caption ISO 200
Figure 14: Qualitative results of our noise translation network with pretrained NAFNet applied to high ISO images from the NIND dataset. Our method effectively removes noise while preserving fine details, producing outputs comparable to low ISO (ISO 200) images. Zoom in for a closer inspection of the visual quality.
@@ 10 Limitations
Figure 15: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet) on images captured by Galaxy S22+ smartphone. Zoom in for a closer inspection of the visual quality.